A silicone diode in a adapter rectifying circuit has a carrier density of {10}^{ - 2} in p material and {10}^{ -22} {m}^{ - 3} in n material, the temperature of the charger changes from 25 degrees Celsius to 40 degrees Celsius , find the change in barrier potential of the diode​
1
Expert's answer
2020-10-12T07:46:36-0400
"\frac{i}{I}=\exp{\frac{eV}{kT}}-1\\\frac{i'}{i}=\frac{\exp{\frac{eV'}{kT'}}-1}{\exp{\frac{eV}{kT}}-1}\\\frac{10^{-22}}{10^{-2}}=\frac{\exp{\frac{(1.6\cdot10^{-19}V'}{(1.38\cdot10^{-23})(273+40)}}-1}{\exp{\frac{(1.6\cdot10^{-19}V}{(1.38\cdot10^{-23})(273+25)}}-1}"

"V'=1.05V\\\frac{\Delta V}{V}=0.05=5\%"